OpenStudy (trwatkins1):

Which equation is related to the given equation? Column A 1. y = 49 + 26 2. 26 = 49 – y 3. 49 – 26 = y 4. –26 = 49 – y Column B A. y – 26 = 49 B. y + 26 = 49

OpenStudy (trwatkins1):

eny idea

OpenStudy (welshfella):

add 26 to both sides of equation A what do you get?

OpenStudy (welshfella):

- equation A in column B

OpenStudy (trwatkins1):

???

OpenStudy (welshfella):

add 26 to both sides of the '=' sign in equation y - 26 = 49

OpenStudy (welshfella):

this will simplify to one of the equations in Column |A

OpenStudy (trwatkins1):

75

OpenStudy (trwatkins1):

i dont understand

OpenStudy (welshfella):

- what equation do you get ?

OpenStudy (welshfella):

y - 26 + 26 = 49 + 26

OpenStudy (welshfella):

now simplify the left side

OpenStudy (trwatkins1):

52-y

OpenStudy (welshfella):

- 26 + 26 does not equal 52

OpenStudy (trwatkins1):

oh that divide thought it was +

OpenStudy (welshfella):

no its a plus

OpenStudy (trwatkins1):

and i feel dumb i forgot negetive

OpenStudy (trwatkins1):

its 0

OpenStudy (welshfella):